# Emergence Learning: A Rising Direction from Emergent Abilities and a Monosemanticity-Based Study

### 17 Dec 2023 v1

@@ 4.1 Setup
Original 80.9 83.2 85.1
MEmeL 81.0 83.4 85.1
@@
Table 3: Results on HKO-7 dataset. We initially train a ConvGRU model for 20k steps to create the base model. Then, we perform finetuning for 8k steps with and without MEmeL, labeled as “MEmeL” and “Original” respectively. The metrics used are B-MSE and B-MAE, where a smaller value indicates better performance. The best results are in bold fonts. The model is fine-tuned three times on each setting for fairness.
@@
Original 312.6$\pm$15.8 997.5$\pm$12.3
MEmeL 311.9$\pm$14.0 995.8$\pm$9.4
To validate our hypothesis on monosemanticity, we conduct experiments on tasks including language, image, and physics simulation to test our MEmeL method:
  • Language Task. Our study is based on the BERT (Pre-training of Deep Bidirectional Transformers) model (Devlin et al., 2019) and the GLUE (General Language Understanding Evaluation) benchmark (Wang et al., 2019). BERT utilizes a transformer architecture and is pretrained on a large corpus of text data. It excels at capturing context and generating high-quality representations of words and sentences. We apply EMemL to the BERT-base model on the GLUE dataset. GLUE serves as a benchmark for natural language understanding tasks, encompassing various tasks such as natural language inference, sentiment analysis, and similarity analysis.
  • Physics Simulation Task. We apply our EMemL to the ConvGRU model on the HKO-7 dataset, which forecasts precipitation based on images of radar echoes. (Shi et al., 2015, 2017). Precipitation forecasting is a challenging task, for both theory-driven and data-driven approaches, as it exhibits complex chaotic dynamics (Xuebin et al., 2017; Wang et al., 2017). ConvGRU is a lightweight module that belongs to the ConvRNN class, a milestone structure that combines CNN and RNN to capture spatiotemporal correlations simultaneously. HKO is a large-scale benchmark dataset from the Hong Kong Observatory (HKO), providing high-resolution radar images spanning multiple years.
We emphasize that our method is a general module applicable to models of any scale. After applying our module, the effectiveness of MEmeL can be evaluated by comparing it with the original model.
All the codes are implemented in PyTorch (Adam et al., 2017), which are available in the supplementary materials. Our experiment is conducted on 4 V100 cards, 8 RTX2080 cards, and 4 RTX3090 cards. To avoid a cheating-style parameter search for improvement, we minimize the tunable hyperparameter during training. For example, after calculating the MS for each layer of neurons, one can introduce various hyperparameters to determine the number of neurons to be deactivated by RD. These hyperparameters include top-$k$, top $p$ percent, and a hard threshold $c$ so that neurons with $MS>c$ are processed, etc. We choose the lightest modification, deactivating the neuron with the top-$1$ MS in each batch, to prove the effectiveness of our methods. More details are given in Appendix. F. Our setting may somehow compromise the improvement in score, but it preserves fairness. See more discussions in Appendix. G.3.
4.2 Main Experiment Result
We present our results on the GLUE dataset in Table. 1. BERTMEmeL is our variant obtained by applying MEmeL to BERTBASE. Our method improves results on 8 tasks and achieves the same results on MRPC. (The results were obtained from our first submission to GLUE. Since our goal is to validate the effectiveness of our method rather than prove its upper bound, we did not invest time in a second submission, although the score of MRPC might be further improved by simple tuning mentioned above.)
The results on ImageNet-1k are given in Table. 2. We finetune the checkpoints provided by Swin-Transformer on three different scales: Tiny (Swin-T), Small (Swin-S), Base (Swin-B). Our method achieves the same score on the Base and Large models, and improves the results on the two smaller versions. A preliminary hypothesis is that image classification might be decomposed into feature detection and estimation of the likelihood of classes based on the presence of these features, which is relatively linear. [Update: By simply increasing the number of neurons to be deactivated from top-1 to top-10, we can outperform Swin-B. Therefore, another reason is that a large model has an great number of neurons in each layer, and deactivating only one for each sample is too mild.]
For the precipitation forecasting task, we first pretrain a ConvGRU model on HKO for 20k steps as the base model. Then, we proceed with training for an additional 8k steps, both with and without MEmeL, for comparison purposes. The results are presented in Table. 3. The metrics used are Balanced Mean Absolute Error (B-MAE) and Balanced Mean Square Error (B-MSE), specifically designed to capture heavy rainfall (Shi et al., 2017) (See Appendix. F). Our method demonstrates better performance compared to the baseline, with lower variation.
5 Related Works
5.1 Artificial Neural Networks and the Increasing Scale
@@
Since S & Walter (1943) first modeled a simple neural network, ANN has been investigated under a scale of several layers in the last century (LeCun et al., 1989; Yann et al., 1995). With improvements in hardware, deeper models can be supported. AlexNet, which uses 8 layers for image classification, achieved excellent performance in the 2012 ImageNet challenge (Krizhevsky et al., 2012). Later on, deeper models such as Inception and VGG increased the scale to tens and hundreds of layers (Szegedy et al., 2015; Simonyan & Zisserman, 2015). The design of critical modules, such as ResNet, also plays a crucial role in ensuring the stability of model training when increasing the depth (He et al., 2016).
@@
After the Transformer was proposed and validated as effective in various areas (Vaswani et al., 2017), both its scale and performance have seen significant growth over the years (Devlin et al., 2019; Liu et al., 2021). In recent years, architectures based on the Transformer have achieved great success at extremely large scales (Floridi & Chiriatti, 2020; Touvron et al., 2023). Investigating the underlying effective properties created by the increase in scale is a highly demanded research direction.
5.2 Mechanistic Interpretability
@@
With the improved performance of neural networks, their black-box nature raises more questions than it answers. To gain a better understanding of and diagnose neural networks, researchers seek mechanistic interpretability (Räuker et al., 2022). They study individual components to understand their functionality and usage, such as neurons for identifying dogs and cars (Chris et al., 2020). As Transformer models demonstrate their superiority in various domains, there is an increasing focus on their interpretability. Geva et al. (2021a) propose that the feed-forward layers of Transformers function as key-value pairs. Geva et al. (2021b) extend this mapping to the embedding space. Although the complexity greatly increases with larger models, the success of these models attracts researchers who strive to find interpretability in the vast sea of neurons (Gurnee et al., 2023; Trenton et al., 2023; Steven et al., 2023a). While the desire to decompose and understand everything is appealing (e.g., achieving monosententicity), it may conflict with the progress of intelligence.
5.3 Debate on the Existence of Emergence
@@
6 Conclusion
In this paper, we discuss the concept of Emergence Learning and its importance for large-scale models. We focus on the decrease of monosemanticity as an important factor for Emergence Learning and propose methods to proactively inhibit monosemantic neurons. Our experiments validate this idea. Additionally, we provide further discussion on the monosemanticity, where researchers prefer monosemantic neurons as they have strong one-to-one correlations and are easy for AI explanation (Gurnee et al., 2023; Dar et al., 2023; Geva et al., 2021a). However, amazing intelligence could be a result of chaos, where monosemantic should be reduced instead. This hypothesis also implies bad news in the safety of AI, where more powerful AI tools (e.g., deep learning models) naturally exhibit bad explainability. In other words, serving as an Uncertainty Principle in the AI area, one may hypothesize that an AI model cannot output both the answer and reason of a complex problem, with perfect accuracy.
Appendix A Examples of Decreasing Monosemanticity in Large Models
Figure 4: Examples of features and their monosemantic neurons from Gurnee et al. (2023). We have taken a screenshot of the distribution of activation values when features present (blue) and not present (orange) in the input context. (a) 70M model: Neurons that activate on tokens in German (left), Italian (middle), and Greek (right), as shown in Figure 15 of Gurnee et al. (2023). (b) 6.9B model: Neurons that activate on tokens in arxiv (left), uspto (middle), and freelaw (right), as shown in Figure 17 of Gurnee et al. (2023).
Even for the neurons that are mostly likely to be monosemantic, larger models show weak reliance on these neurons when given corresponding inputs. In Section 5.2 of the Gurnee et al. (2023), the authors investigate neurons that serve as context features, which are expected to have the most monosemanticity. However, deactivating these neurons only results in a small decrease in performance for larger models. For example, ablating a single neuron for “is_french” in a 70M parameter model leads to an average loss increase of 8% per French sequence, while the increase in loss is only 0.2% in the 6.9B model.
Moreover, when examining the activation values, neurons in smaller models show more variation when the corresponding feature is present or absent in the input. Figure. 4 presents figures from Gurnee et al. (2023). The gaps between peaks of value when features are in (blue) and not in (orange) inputs are much larger in 70M model (Figure. 4(a)) than in 6.9B model (Figure. 4(b)), indicating that the sensitivity of neuron on a single feature is much more significant in smaller scales, thus more monosemantic.
When creating interpretable key-value pairs for neurons, it is statistically uncommon to find a one-to-one correlation of monosemanticity in large models. For example, in a 16-layer transformer language model, Geva et al. (2021b) find an average of 3.6 expert identifiable features per neuron.
Appendix B Proof of Lemma. 3.2
@@ Appendix A Examples of Decreasing Monosemanticity in Large Models
$\mu_{m}=\frac{\sum_{i=1}^{m}h_{(i)}}{m},\mu_{b}^{\prime}=\frac{\sum_{i=1}^{b}h_{(m+i)}}{b}$
@@
$\mu_{m+b}$ $=\frac{\sum_{i=1}^{m+b}h_{(i)}}{m+b}$ (9)
$=\frac{\sum_{i=1}^{m}h_{(i)}+\sum_{i=1}^{b}h_{(m+i)}}{m+b}$ (10)
$=\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b}$ (11)
@@
$\upsilon_{m}=\frac{\sum_{i=1}^{m}(h_{(i)}-\mu_{m})^{2}}{m-1},\upsilon^{\prime}_{b}=\frac{\sum_{i=1}^{b}(h_{(m+i)}-\mu^{\prime}_{b})^{2}}{b}.$
@@
$\upsilon_{m+b}=$ $\frac{\sum_{i=1}^{m+b}(h_{(i)}-\mu_{m+b})^{2}}{m+b-1}=\frac{\sum_{i=1}^{m+b}(h_{(i)}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}}{m+b-1}$ (12)
$=$ $\frac{\sum_{i=1}^{b}(h_{(m+i)}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}+\sum_{i=1}^{m}(h_{(i)}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}}{m+b-1}$ (13)
$=$ $\frac{\sum_{i=1}^{b}\left[h_{(m+i)}-\mu^{\prime}_{b}-\frac{m(\mu_{m}-\mu^{\prime}_{b})}{m+b}\right]^{2}+\sum_{i=1}^{m}\left[h_{(i)}-\mu_{m}+\frac{b(\mu_{m}-\mu^{\prime}_{b})}{m+b}\right]^{2}}{m+b-1}$ (14)
$=$ $\frac{b\left(\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{b}\left(h_{(i+m)}-\mu^{\prime}_{b}\right)^{2}-\sum_{i=1}^{b}\left[2(h_{(m+i)}-\mu^{\prime}_{b})\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$ (15)
$+\frac{m\left(\frac{b(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{m}\left(h_{(i)}-\mu_{m}\right)^{2}+\sum_{i=1}^{m}\left[2(h_{(i)}-\mu_{m})\frac{b(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$ (16)
$=$ $\frac{mb^{2}(\mu_{m}-\mu_{b}^{\prime})^{2}+bm^{2}(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)^{2}}+\frac{\sum_{i=1}^{b}\left(h_{(m+i)}-\mu_{b}^{\prime}\right)^{2}-0+(m-1)\upsilon_{m}+0}{m+b-1}$ (17)
$=$ $\frac{mb(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)}+\frac{b\upsilon^{\prime}_{b}+(m-1)\upsilon_{m}}{m+b-1},$ (18)
where the third terms in Equation. (15) and Equation. (16) equal to zero as $\sum_{i=1}^{b}(h_{(m+i)}-\mu^{\prime}_{b})=\sum_{i=1}^{b}(h_{(m+i)})-b\mu^{\prime}_{b}=0,\sum_{i=1}^{m}(h_{(i)}-\mu_{m})=\sum_{i=1}^{m}(h_{(i)})-m\mu_{m}=0$ and other terms are constant factors.
Appendix C Proof and Discussion of Lemma. 3.3
@@
Recall that based on Reversed Deactivation, our gradient flow originates from the following equations
$Z^{\prime}=$ $f_{2}(H^{\prime},X)=f_{2}([-h+(h+\bar{h})_{ng},H^{-},X]),$ (19)
$H=$ $[h,H^{-}]=f_{1}(X),$ (20)
where $\cdot_{ng}$ denotes “no gradient” scalar and ${H^{-}}\subset H$ are the neurons in $H$ except $h$.
Though a bit of abusing the symbol, to give a clearer representation for gradient update, we use $h$ to refer to the neuron, $h_{a}$ refers to its activation value which leads to an ideal $Z$. $h_{o}=h_{a}$ as the value of $h$ before processed by RD, $h_{u}$ for the value updated after gradient descent. We still use $\bar{h}$ as the mean value of $h$ and represent the difference between $h$ and $\bar{h}$ as an updating variable $t=h-\bar{h}$. Note that except $h$ and $t$, other definitions in this paragraph are untrainable scalars.
Here we focusing on $h-$related parts of Equation. (19) and (20). We denote parameters of $f_{1}$ as $\theta$ and rewrite Equation. (20) for gradient calculation and parameter update:
$h=$ $g(\theta)+\bar{h}$ (21)
$\Rightarrow t=$ $g(\theta),$ (22)
where $g$ is $h$-related functions of $f_{1}$.
Here we fix other states and focus on the interaction between $h$, $\theta$, and loss $\mathcal{L}(Z^{\prime})$. We can calculate the gradient for $\theta$:
$\nabla\mathcal{L}(\theta)=$ $\frac{\partial{\mathcal{L}(Z^{\prime})}}{\partial\theta}$ (23)
$=$ $\frac{\partial{\mathcal{L}(f_{2}(H^{\prime}))}}{\partial t}\frac{\partial t}{\partial\theta}$ (24)
$=$ $\frac{\partial{\mathcal{L}(f_{2})}}{\partial(h^{\prime}-h_{a})}\frac{\partial(h^{\prime}-h_{a})}{\partial t}\frac{\partial t}{\partial\theta}$ (25)
$=$ $-\nabla f(\delta)\nabla g(\theta),$ (26)
where $\frac{\partial t}{\partial\theta}=\nabla g(\theta)$ as the gradient of Equation. (22) and $\frac{\partial(h^{\prime}-h_{a})}{\partial t}=\frac{\partial(-h+h_{a}+\bar{h}-h_{a})}{\partial(h-\bar{h})}=-1$. $\nabla f(\delta)$ denotes $\frac{\partial{\mathcal{L}(f_{2})}}{\partial(h^{\prime}-h_{a})}.$ According to our assumption that $f(\delta)$ the monotonically increases with $|h^{\prime}-h_{a}|$, $\nabla f(\delta)\geq 0$ when $h^{\prime}\geq h_{a}$ so that $h^{\prime}-h_{a}=|h^{\prime}-h_{a}|$, which means $\bar{h}\geq h_{a}$ and $t\leq 0$. Also, we can derive that $t\geq 0$ when $\nabla f(\delta)\leq 0$.

### 15 Feb 2024 v2

@@ 4.1. Experimental Setup
Original 80.9 83.2 85.1 1003.41 309.96
Naive (a) 81.0 83.4 84.6 1003.56 309.83
Naive (b) 81.0 83.4 85.1 1003.40 310.13
MEmeL 81.1 83.4 85.1 1003.25 209.94
MEmeL-Tune 81.1 83.5 85.2 998.81 298.16
Table 3. Validation experiments conducted on the Swin-B model. We record the Decrease Ratios and Update Scales of 10k neurons. The model that utilizes our Reverse Deactivation method is compared with those using two Naive methods and the original Swin-B.
Methods Original Naive (a) Naive (b) Reverse Deactivation
Average Decrease Ratio 0.003% -0.017% -0.044% 0.013%
Average Total Update Ratio 0.052% 0.118% 0.161% 0.189%
4.1.1. Data Sets, Base Models, and Tasks
To validate the conjecture of monosemantic neurons, our inhibition method, and corresponding theories, we apply MEmel to milestone models such as Bert, Transformer, and ConvRNN on various tasks (e.g., language, image, and physics simulation), respectively.
  • Language Task. We apply MEmeL to the BERT (Pre-training of Deep Bidirectional Transformers) model (Devlin et al., 2019) on the GLUE (General Language Understanding Evaluation) benchmark (Wang et al., 2019). BERT utilizes a transformer architecture and is pretrained on a large corpus of text data. It excels at capturing context and generating high-quality representations of words and sentences. GLUE serves as a benchmark for natural language understanding tasks, encompassing various tasks such as natural language inference, sentiment analysis, and similarity analysis.
  • Physics Simulation Task. We apply our MEmeL to the ConvGRU model on the HKO-7 dataset (Shi et al., 2017), which forecasts precipitation based on images of radar echoes (Shi et al., 2015). ConvGRU is a lightweight module that belongs to the ConvRNN class, a milestone structure that combines CNN and RNN to capture spatiotemporal correlations simultaneously. HKO is a large-scale benchmark dataset from the Hong Kong Observatory (HKO), providing high-resolution radar images spanning multiple years. Generally, precipitation forecasting is a challenging task, for both theory-driven and data-driven approaches, as it exhibits complex chaotic dynamics (Xuebin et al., 2017; Wang et al., 2017).
The configuration of above base models generally follow their original settings (see more details in Appx. E.1). The metrics can be found in Appx. E.2. We emphasize that our method is a general module applicable to models of any scale. After applying our module, the effectiveness of MEmeL can be evaluated by comparing it with the original model.
4.1.2. Hyper-parameter Setting
This paper focuses on the performance improvement that is brought by introducing reverse deactivation into milestone models. To fairly validate the effectiveness of the proposed MEmel, we deactivate the neuron with only top-$1$ MS in each batch (recorded as $\rm{MEmeL}$). In other words, we give up the way of determining the number of neurons to be deactivated through hyper-parameter tuning. Therefore, setting may somehow compromise the score improvement, but it preserves fairness. Besides, we also display the results with tuned hyper-parameters (recorded as $\rm{MEmeL-Tune}$). See more details in Appx. E.3.
4.1.3. Implementation
All the codes are implemented in PyTorch (Adam et al., 2017), which is available in the supplementary materials. Our experiment is conducted on 4 V100 GPU cards. Our code are available through the anonymous repo.
4.2. Main Experiment Result
Tab. 1 and Tab. 2 demonstrate the performance of MEmeL incorporated with BERT, Swin-Transformer, and ConvGRU, respectively. The “Original” method denotes the raw model, “Naive (a)” and “Naive (b)” are corresponding to the original model incorporated with naive inhibition methods (see details in Eq. (10)). By comparing the original method and the method enhanced with MEmeL, we can see that the MEmeL (especially MEmeL-Tune) achieves better or comparable results on different tasks, different data sets, and different base models. This validates that neural networks can achieve better results by deactivating monosemantic neurons. By comparing Naive (a), (b) with MEmeL (and MEmeL-Tune), we can easily observe that naive deactivation methods may intensify the monosemanticity of neural networks, which may lead to inferior performance improvement as discussed in Sec. 3.2.1.
4.3. Case Study on Deactivating Monosemantic Neurons
To validate that the monosemanticity is indeed inhibited by our reverse deactivation, we conducted experiments on the ImageNet-1k dataset using the Swin-Transformer model as shown in Tab. 3. We collected 10k samples for 4 different settings, where the modification of $h$ was done using Original (no modification), and two naive methods (a) and (b), and Reverse Deactivation in Sec. 3.2.2.
We feed inputs $\mathbf{x}$ to the model again to check how $z$ is optimized after updating the model with $\mathbf{x}$. The new value is denoted as $z^{\prime}$. Then, the Decrease Ratio represents how the monosemanticity is decreased upon $z$: Decrease $\textit{Decrease Ratio}=\left(1-\phi(z^{\prime})/\phi(z)\right)\times 100\%$. Without any modification, monosemanticity showed a slight decrease with a small positive average decrease ratio (0.003%) in the original model. Naive methods (a) and (b) intensify monosemanticity since their decrease ratios are negative, which is consistent with our analysis in Sec. 3.2.1. On the contrary, reverse deactivation has the most significant impact on decreasing monosemanticity (0.013%). But the value is relatively small. That is mainly because the small learning rate ($2\times 10^{-6}$) usually leads to a small update step in training. To provide a clear illustration, we further define the Total Update Ratio as Total Update $\textit{Total Update Ratio}=\left|z^{\prime}/z-1\right|\times 100\%$. The scale of modification on $z$ and on $\phi(\cdot)$ is compatible (e.g., 0.013% versus 0.189% for RD). Thus, the small scale of the Decrease Ratio is due to the small learning rate, indicating a stable impact on the inhibition of monosemanticity using our method.
4.4. Potential and Limitation of MEmeL
According to our hypothesis, MEmeL induces the model to accumulate general and abstract functionality instead of monosemanticity for a specific task, which is consistent with the goal of per-taining. Although MEmeL achieves good results during fine-tuning (demonstrated at Main Experiments in Sec. 4.2), the improvement is expected to be even greater when it is applied to the pre-training phase.
The main obstacle is the high computational resource cost required for per-taining. Currently, we have only completed validation on relatively small precipitation forecasting datasets. In Tab. 4, the model pre-trained with MEmeL (P-MEmeL) outperforms the one without it (P-Original). Based on these two models, the same finetuning process is conducted without MEmeL. Using MEmeL during pretraining improves B-MAE by 0.18% and B-MSE by 1.29% (T-MEmeL). In contrast, using MEmeL during finetuning improves B-MAE by 0.02% B-MSE by 0.01% (MEmeL in Table. 2). The results validate our hypothesis.
@@ 4.4. Potential and Limitation of MEmeL
Table 4. Validation experiments conducted on the HKO-7 dataset. In addition, we pretrain a ConvGRU model for 20k steps using MEmeL, labeled as “P-MEmeL”. The “P-Original” model is pretrained based on the original model and is used in the main experiment. Based on these two models, we perform finetuning for 2k steps using the original model, labeled as “T-MEmeL” and “T-Original”, respectively. The metrics used for evaluation were B-MSE and B-MAE, where a smaller value indicates better performance. The best results are shown in bold fonts.
@@
P-Original 1004.25 311.54
P-MEmeL 1000.98 306.67
T-Original 1003.41 309.96
T-MEmeL 1001.56 305.96
5. Conclusion
Different from the literature, we hypothesize a key factor that highly promote the performance of large neural networks: the reduction of monosemantic neurons. There is no unified quantitative evaluation metric and simply banning monosemantic neurons does not promote polysemanticity in neural networks. Therefore, we propose to learn from emergence and present a study on proactively inhibiting the monosemantic neurons in this paper. More specifically, we first propose a new metric to measure the monosemanticity of neurons with the guarantee of efficiency for online computation, then introduce a theoretically supported method to suppress monosemantic neurons and proactively promote the ratios of polysemantic neurons in training neural networks. We validate our conjecture that monosemanticity brings about performance change at different model scales on a variety of neural networks and benchmark datasets in different areas, including language, image, and physics simulation tasks. Further experiments validate our analysis and theory regarding the inhibition of monosemanticity.
Unfortunately, extending this research to very large data sets or models (e.g., large language models) is appealing yet impossible for research departments due to limited resources. Therefore, we are trying to find ways to extend this paper to extremely large models trained on large data sets as future work.
Appendix A An Example Analogy that Helps Illustrate Our Motivation
We display an example with neuroscience analogy to better illustrate our motivation. For more relationships with neuroscience, see discussions in subsection. G.1.
From the perspective of scale differences, when the model is small, the limited number of neurons makes it difficult to abstract and analyze the input, and to provide functional support for integrating a large amount of scattered information (Figure. 4(a)). To achieve good results, neurons tend to specialize in certain features, resulting in stronger monosemanticity (Figure. 4(b)).
Similarly, we can analogize large models to adults, allowing them to learn and master solving skills related to integration (Figure. 4(c)). This is because adults have accumulated knowledge and have a more mature brain development, similar to the huge training volume and parameter size of large models. In this process, adults gradually reduce the reliance on rote memorization and improve their reasoning abilities for new problems (Figure. 4(d)). This can explain why large models have fewer monosemantic neurons and stronger generalization abilities in complex tasks.
Appendix B Related Works
B.1. Artificial Neural Networks and the Increasing Scale
@@
Since S and Walter (1943) first modeled a simple neural network, ANN has been investigated under a scale of several layers in the last century (LeCun et al., 1989; Yann et al., 1995). With improvements in hardware, deeper models can be supported. AlexNet, which uses 8 layers for image classification, achieved excellent performance in the 2012 ImageNet challenge (Krizhevsky et al., 2012). Later on, deeper models such as Inception and VGG increased the scale to tens and hundreds of layers (Szegedy et al., 2015; Simonyan and Zisserman, 2015). The design of critical modules, such as ResNet, also plays a crucial role in ensuring the stability of model training when increasing the depth (He et al., 2016).
@@
After the Transformer was proposed and validated as effective in various areas (Vaswani et al., 2017), both its scale and performance have seen significant growth over the years (Devlin et al., 2019; Liu et al., 2021). In recent years, architectures based on the Transformer have achieved great success at extremely large scales (Floridi and Chiriatti, 2020; Touvron et al., 2023). Investigating the underlying effective properties created by the increase in scale is a highly demanded research direction.
B.2. Mechanistic Interpretability
@@
With the improved performance of neural networks, their black-box nature raises more questions than it answers. To gain a better understanding of and diagnose neural networks, researchers seek mechanistic interpretability (Räuker et al., 2022). They study individual components to understand their functionality and usage, such as neurons for identifying dogs and cars (Chris et al., 2020). As Transformer models demonstrate their superiority in various domains, there is an increasing focus on their interpretability. Geva et al. (2021) propose that the feed-forward layers of Transformers function as key-value pairs. Dar et al. (2023) extend this mapping to the embedding space. Although the complexity greatly increases with larger models, the success of these models attracts researchers who strive to find interpretability in the vast sea of neurons (Gurnee et al., 2023; Trenton et al., 2023; Steven et al., 2023b). In their work, Elhage et al. (2022) introduce the softmax linear unit to create monosemantic models. Additionally, Gurnee et al. (2023) modify the sparse probe by using multiple heads to classify neurons with decreasing monosemanticity. Trenton et al. (2023) construct more powerful and complex probes to construct mappings between neurons and features. While the desire to decompose and understand everything is appealing (e.g., achieving monosententicity), it may conflict with the progress of intelligence.
B.3. Debate on the Existence of Emergence
@@
Appendix C Proof of Sec. 3
C.1. Proof of Lemma. 3.2
@@
$\mu_{m}=\frac{\sum_{i=1}^{m}z^{[i]}}{m},\mu_{b}^{\prime}=\frac{\sum_{i=1}^{b}z^{[m+i]}}{b}.$
@@
(15) $\mu_{m+b}$ $=\frac{\sum_{i=1}^{m+b}z^{[i]}}{m+b}$
(16) $=\frac{\sum_{i=1}^{m}z^{[i]}+\sum_{i=1}^{b}z^{[m+i]}}{m+b}$
(17) $=\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b}.$
@@
$\upsilon_{m}=\frac{\sum_{i=1}^{m}(z^{[i]}-\mu_{m})^{2}}{m-1},\upsilon^{\prime}_{b}=\frac{\sum_{i=1}^{b}(z^{[m+i]}-\mu^{\prime}_{b})^{2}}{b}.$
@@
(18) $\upsilon_{m+b}=$ $\frac{\sum_{i=1}^{m+b}(z^{[i]}-\mu_{m+b})^{2}}{m+b-1}=\frac{\sum_{i=1}^{m+b}(z^{[i]}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}}{m+b-1}$
(19) $=$ $\frac{\sum_{i=1}^{b}(z^{[m+i]}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}+\sum_{i=1}^{m}(z^{[i]}-\frac{m\mu_{m}+b\mu^{\prime}_{b}}{m+b})^{2}}{m+b-1}$
(20) $=$ $\frac{\sum_{i=1}^{b}\left[z^{[m+i]}-\mu^{\prime}_{b}-\frac{m(\mu_{m}-\mu^{\prime}_{b})}{m+b}\right]^{2}+\sum_{i=1}^{m}\left[z^{[i]}-\mu_{m}+\frac{b(\mu_{m}-\mu^{\prime}_{b})}{m+b}\right]^{2}}{m+b-1}$
(21) $=$ $\frac{b\left(\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{b}\left(h_{(i+m)}-\mu^{\prime}_{b}\right)^{2}-\sum_{i=1}^{b}\left[2(z^{[m+i]}-\mu^{\prime}_{b})\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$
(22) $+\frac{m\left(\frac{b(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{m}\left(z^{[i]}-\mu_{m}\right)^{2}+\sum_{i=1}^{m}\left[2(z^{[i]}-\mu_{m})\frac{b(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$
(23) $=$ $\frac{mb^{2}(\mu_{m}-\mu_{b}^{\prime})^{2}+bm^{2}(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)^{2}}+\frac{\sum_{i=1}^{b}\left(z^{[m+i]}-\mu_{b}^{\prime}\right)^{2}-0+(m-1)\upsilon_{m}+0}{m+b-1}$
(24) $=$ $\frac{mb(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)}+\frac{b\upsilon^{\prime}_{b}+(m-1)\upsilon_{m}}{m+b-1},$
where the third terms in Equation. (21) and Equation. (22) equal to zero as $\sum_{i=1}^{b}(z^{[m+i]}-\mu^{\prime}_{b})=\sum_{i=1}^{b}(z^{[m+i]})-b\mu^{\prime}_{b}=0,\sum_{i=1}^{m}(z^{[i]}-\mu_{m})=\sum_{i=1}^{m}(z^{[i]})-m\mu_{m}=0$ and other terms are constant factors. ∎
C.2. Proof and Discussion of Lemma. 3.3
@@
Recall that based on Reversed Deactivation, our gradient flow originates from the following equations:
(25) $\mathbf{o}^{\prime}=$ $f_{2}(\mathbf{z}^{\prime},\mathbf{x})=f_{2}([-z+(z+\bar{z})_{ng},\mathbf{z}^{-},\mathbf{x}]),$
(26) $\mathbf{z}=$ $[z,\mathbf{z}^{-}]=f_{1}(\mathbf{x}),$
where $\cdot_{ng}$ denotes “no gradient” scalar and ${\mathbf{z}^{-}}\subset\mathbf{z}$ are the neurons in $\mathbf{z}$ except $z$.
Though a bit of abusing the symbol, to give a clearer representation for gradient update, we use $z$ to refer to the neuron, $z_{a}$ refers to its activation value which leads to an ideal $\mathbf{o}$. $z_{o}=z_{a}$ as the value of $z$ before processed by RD, $z^{*}$ for the value updated after gradient descent. We still use $\bar{z}$ as the mean value of $z$ and represent the difference between $z$ and $\bar{z}$ as an updating variable $t=z-\bar{z}$. Note that except $z$ and $t$, other definitions in this paragraph are untrainable scalars.
Here we focusing on $z-$related parts of Equation. (25) and (26). We denote parameters of $f_{1}$ as $\theta\in\mathbf{W}$ and rewrite Equation. (26) for gradient calculation and parameter update:
(27) $z=$ $g(\theta)+\bar{h}$
(28) $\Rightarrow t=$ $g(\theta),$
where $g$ is $z$-related functions of $f_{1}$.
Here we fix other states and focus on the interaction between $z$, $\theta$, and loss $\mathcal{L}(\mathbf{o}^{\prime})$. We can calculate the gradient for $\theta$:
(29) $\nabla\mathcal{L}(\theta)=$ $\frac{\partial{\mathcal{L}(\mathbf{o}^{\prime})}}{\partial\theta}$
(30) $=$ $\frac{\partial{\mathcal{L}(f_{2}(\mathbf{z}^{\prime}))}}{\partial t}\frac{\partial t}{\partial\theta}$
(31) $=$ $\frac{\partial{\mathcal{L}(f_{2}(\mathbf{z}^{\prime}))}}{\partial(z^{\prime}-z_{a})}\frac{\partial(z^{\prime}-z_{a})}{\partial t}\frac{\partial t}{\partial\theta}$
(32) $=$ $-\nabla f(\delta)\nabla g(\theta),$
where $\frac{\partial t}{\partial\theta}=\nabla g(\theta)$ as the gradient of Equation. (28) and $\frac{\partial(z^{\prime}-z_{a})}{\partial t}=\frac{\partial(-z+z_{a}+\bar{z}-z_{a})}{\partial(z-\bar{z})}=-1$. $\nabla f(\delta)$ denotes $\frac{\partial{\mathcal{L}(f_{2}(\mathbf{z}^{\prime}))}}{\partial(z^{\prime}-z_{a})}.$ According to our assumption that $f(\delta)$ the monotonically increases with $|z^{\prime}-z_{a}|$, $\nabla f(\delta)\geq 0$ when $z^{\prime}\geq z_{a}$ so that $z^{\prime}-z_{a}=|z^{\prime}-z_{a}|$, which means $\bar{z}\geq z_{a}$ and $t\leq 0$. Also, we can derive that $t\geq 0$ when $\nabla f(\delta)\leq 0$.